Express the higher-order differential equation as a matrix system in normal form. (1-1)/ - 2ty + 2y = 0 (Legendres equatio
Answer #1

The solution to the above problem is given below. Please review the answer and let me know if you have any doubts.

See (I-4) - 2ty + 19 = 0 let =2 -0 24=y 3 2 =y 222 ㅋ 일=y 키 길=y Soo (1-t) - 21 2, + 2244 = 0 (1-t) ~ ? Rt 샤 로 4 (2) fro

Hence D is the correct option.